OpenStudy (jazzyfa30):

please help Exponential and Logarithmic Equations solve each equation. Give an exact solution. log(base4)^(3x-2)=2

OpenStudy (jazzyfa30):

@ganeshie8

OpenStudy (mathmale):

\[\log(base4)^(3x-2)=2\rightarrow \log_{4}(3x-2)=2 \]

OpenStudy (mathmale):

It's important that you know and understand the concept of inverse functions as you start solving this particular problem. Are you aware that \[\log 10^x = x? ~~~\ln e^5=5? If~so, ~what~is~4^{\log_{4} 2}?\]

OpenStudy (mathmale):